### v1

@@ III.2 Predictions of the differential cross sections for $K^{-}p\to K\Xi(1530)$
In addition to the cross sections, the differential cross sections for the involved processes depending on $\cos\theta$ are predicted with the fitted parameters in the present work, where $\theta$ is the angle between the outgoing $\Xi(1530)$ and the ingoing kaon beam direction. The differential cross sections with several typical center-of-mass energies for $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$ are presented in Figs. 9 and  10, respectively. Here, we take the center-of-mass energy $\sqrt{s}$ as 2.05, 2.10, 2.15, 2.20, 2.25, 2.30 GeV, respectively. From the figure one can notice that the differential cross sections reach the maximum at the backward angle limit when $\sqrt{s}=2.10$ and 2.15 GeV, while when $\sqrt{s}=2.25$ GeV, the differential cross sections reach its maximum at the forward angle limit. In addition, by comparing Model A and Model B, one can find that the differential cross sections of Model B are generally greater than that of Model A, which are consistent with the relevant cross sections. Moreover, the differential cross sections for $K^{-}p\rightarrow K^{0}\Xi(1530)^{0}$ are presented in Fig. 10. Similarly, the black solid curves and black dotted curves correspond to Model A and Model B, respectively. In most cases, the differential cross sections are similar for $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$, since the contributions from the $\Sigma$ hyperon are the same for both processes, and those from above threshold $\Lambda$ hyperon are also similar. These differential cross sections predicted in the present work could be crucial tests to the present estimations, which should be accessible by the kaon experiments at J-PARC.
@@ IV SUMMARY
The ground $\Xi$ hyperon state, $\Xi(1314)$ has been comprehensively studied over the years, and the researches on its productions in the $K^{-}p$ scattering processes have been performed in the literature. However, the researches on the productions of $\Xi(1530)$ are limited due to the scarcity of the experimental data. In the present work, we investigate the production of $\Xi(1530)^{0}$ and $\Xi(1530)^{-}$ in the $K^{-}p$ scattering process simultaneously, where nine $\Lambda/\Sigma$ hyperons and their resonances are involved in the $s$- and $u$-channel processes. Checking the experimental data, we find that the measured cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ are inconsistent with each other in the range $\sqrt{s}=[2.087,2.168]\ \mathrm{GeV}$. Thus, in the present work, we take two different strategies. In Model A, the weight of the experimental data are the same, while in Model B, the experimental data have different weight. Based on a comparative analysis of the fit results in two models, we conclude that Model A demonstrates a better global agreement with the experimental data than Model B.
In addition to the cross sections, the individual contributions from different intermediate states to the cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$ are also estimated with the fitted parameters. Our results indicate that the cross section resulted from $\Sigma(1193)$ intermediate process is predominant. Moreover, the different cross sections are predicted for both the charged and neutral processes with several typical center-of-mass energies, which could be tested by further experimental measurements at J-PARC in the future.
@@ ACKNOWLEDGMENTS
This work is partly supported by the National Natural Science Foundation of China under the Grant Nos. 12175037 and 12335001, as well as supported, in part, by National Key Research and Development Program under the contract No. 2024YFA1610503. PW acknowledges support from Natural Science Foundation of Jiangsu Province (Grant No. BK20210201), Fundamental Research Funds for the Central Universities, Excellent Scholar Project of Southeast University (Class A), and the Big Data Computing Center of Southeast University.

### v2

@@ III.2 Predictions of the differential cross sections for $K^{-}p\to K\Xi(1530)$
In addition to the cross sections, the differential cross sections for the involved processes depending on $\cos\theta$ are predicted with the fitted parameters in the present work, where $\theta$ is the angle between the outgoing $\Xi(1530)$ and the ingoing kaon beam direction. The differential cross sections with several typical center-of-mass energies for $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$ are presented in Figs. 9 and  10, respectively. Here, we take the center-of-mass energy $\sqrt{s}$ as 2.05, 2.10, 2.15, 2.20, 2.25, and 2.30 GeV, respectively. From the figures one can notice that the differential cross sections reach the maximum at the backward angle limit when $\sqrt{s}=2.10$ and 2.15 GeV, while when $\sqrt{s}=2.25$ GeV, the differential cross sections reach the maximum at the forward angle limit. In addition, by comparing model A and model B, one can find that the differential cross sections of model B are generally greater than that of model A, which are consistent with the relevant cross sections. Moreover, the differential cross sections for $K^{-}p\rightarrow K^{0}\Xi(1530)^{0}$ are presented in Fig. 10. Similarly, the black solid curves and black dotted curves correspond to model A and model B, respectively. In most cases, the differential cross sections are similar for $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$, since the contributions from the $\Sigma$ hyperon are the same for both processes, and those from the above-threshold $\Lambda$ hyperon are also similar. These differential cross sections predicted in the present work could be crucial tests for the present estimations, which should be accessible by the kaon experiments at J-PARC.
@@ IV SUMMARY
The ground $\Xi$ hyperon state, $\Xi(1314)$ has been comprehensively studied over the past years, and the researches on its productions in the $K^{-}p$ scattering processes have been performed in the literature. However, the research on the productions of $\Xi(1530)$ is limited due to the scarcity of the experimental data. In the present work, we investigate the production of $\Xi(1530)^{0}$ and $\Xi(1530)^{-}$ in the $K^{-}p$ scattering process simultaneously, where nine $\Lambda$ and $\Sigma$ hyperons and their resonances are involved in the $s$- and $u$-channel processes. Checking the experimental data, we find that the measured cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ are inconsistent with each other in the range $\sqrt{s}=[2.087,2.168]\ \mathrm{GeV}$. Thus, in the present work, we take two different fitting strategies. In model A, the weights of the experimental data are the same, while in model B, the experimental data have different weights. Based on a comparative analysis of the fit results in two models, we conclude that model A demonstrates a better global agreement with the experimental data than model B.
In addition to the cross sections, the individual contributions from different intermediate states to the cross sections for $K^{-}p\to K^{+}\Xi(1530)^{-}$ and $K^{-}p\to K^{0}\Xi(1530)^{0}$ are also estimated with the fitted parameters. Our results indicate that the cross section resulted from the $\Sigma(1193)$ intermediate process is predominant. Moreover, the different cross sections are predicted for both the charged and neutral processes with several typical center-of-mass energies, which could be tested by further experimental measurements at J-PARC in the future.
@@ ACKNOWLEDGMENTS
This work is partly supported by the National Natural Science Foundation of China under Grants No. 12175037 and No. 12335001, as well as supported, in part, by the National Key Research and Development Program under Contract No. 2024YFA1610503. PW acknowledges support from the Natural Science Foundation of Jiangsu Province (Grant No. BK20210201), Fundamental Research Funds for the Central Universities, Excellent Scholar Project of Southeast University (Class A), and the Big Data Computing Center of Southeast University.